OpenStudy (anonymous):

5/3 divided by 7/3 as a fraction

OpenStudy (astrophysics):

Hint: \[\huge \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } \implies \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

Thanks

OpenStudy (eric_d):

Welcome to Open Study !!! @SavanahLewis

OpenStudy (anonymous):

Thank u @eric_d

OpenStudy (astrophysics):

What did you get as your answer?

OpenStudy (anonymous):